# Cross-Modality Embedding of Force and Language for Natural Human-Robot Communication

### v1

@@ Abstract
00footnotetext: Detailed therapy sessions can be seen in our online appendix: https://shared-language-force-embedding.github.io/therapy-sessions/
@@ I Introduction
As stipulated in Asimov’s laws, robots must perform tasks based on human instruction. A central challenge in robotics has been developing effective ways for humans to communicate and give instructions to robots. Early roboticists attempted to develop robot languages, e.g. VAL [19], to describe desired robot actions and tasks. However, despite some success, it became apparent that certain motions and behaviors are difficult, inefficient, or impossible to describe through language alone. This limitation is particularly evident in tasks requiring environmental contact and force/compliance control, where forces and moments are not directly visible. Such manipulative tasks, now termed contact-rich manipulation [20, 4], often involve subconscious knowledge that humans find difficult to articulate through any form of language. To address this challenge, roboticists developed alternative approaches that bypass the need to translate subconscious skills into language. These approaches include teaching by demonstration [3], programming by demonstration [7], skill acquisition [17] and, more recently, imitation learning [26].
Figure 1: Physical therapist from Spaulding Rehabilitation Hospital is seen demonstrating ’hamstring curl’ therapy on the patient with neurological injuries 22footnotemark: 2. She is instructing a patient to move ’gently forward’ while providing an assistive force.
@@
    A framework for cross-modality embedding of force profiles and words, enabling translation between physical force curves and natural language descriptions;
@@ III-C1 Minimal Viable Vocabulary
III-C2 Extended GloVe Vocabulary
GloVe Embeddings: We make use of the GloVe (Global Vectors for Word Representation) word embeddings [21], a pretrained model containing 20,000 words where each word is mapped to a 50-dimensional vector that captures semantic relationships. For phrases of up to three words, we concatenate the embeddings (with padding if needed) to form a $150$-dimensional phrase vector. This continuous representation encodes rich semantic relationships and supports similarity-based operations within the embedding space.
Figure 3: Conceptual illustration of the desired properties of the cross-modality latent space. A pair of corresponding force profile and phrase should be located near each other, measured by a distance metric like Euclidean distance or cosine similarity. However, force profiles and phrases that do not correspond should be positioned far away. This demonstrates that similar inputs would be close together and dissimilar inputs would be far apart in the latent space.
@@ III-D Cross-Modality Embedding (Shared Latent Space)
Figure 5: Architecture diagram of the cross-modality dual autoencoder that represents phrases as a concatenation of three $50\times 1$ GloVe ([21]) word embeddings. The phrase embedding input is then passed into the phrase autoencoder, which encodes it into the shared latent space. There the $16\times 1$ embedding can take 2 paths: either be decoded back into a phrase or be translated into a force profile by using the force profile decoder. Force profile inputs are first preprocessed to extract meaningful features for the force profile autoencoder to digest. They are then encoded into the shared latent space. Like phrase inputs, they can either be decoded back into a force profile or be translated into a phrase by using the phrase decoder.
@@ V-A2 Phrase Input
Since neural networks cannot directly process raw text, each phrase must be converted into a fixed-size vector representation that preserves its semantics. We consider one of the two embedding approaches: Binary Phrase Vectors from the Minimal Viable Vocabulary, or GloVe embeddings from Extended Glove Vocabulary (Section  III-C).
@@ V-A3 Dual Autoencoders
Our framework (see Fig. 5) employs two autoencoders [15] —one for force profiles and one for phrases—to map inputs from distinct modalities into a shared latent space. Each autoencoder consists of an encoder network that compresses its input into a fixed-size latent vector (the “bottleneck” representation) and a decoder network that reconstructs the original input from this latent representation.
@@ V-B Multitask Learning
$\mathcal{L}=k_{r}\mathcal{L}_{r}+k_{z}\mathcal{L}_{z}+k_{t}\mathcal{L}_{t},$ (5)
@@ V-B1 Reconstruction Loss ($\mathcal{L}_{r}$)
It measures how accurately each autoencoder (force and phrase) reproduces its own input from the latent vector. For force profiles, we use mean squared error; for phrases, the reconstruction metric depends on the chosen representation, with cross-entropy for binary phrase vectors and mean squared error for GloVe embeddings.
@@ V-C Data Collection
To train and evaluate the cross force-language embedding, we collected data from 10 participants using a UR robotic manipulator. The study was approved by the Institutional Review Board (IRB Protocol #2212000845R001) to ensure ethical human subject research guidelines were followed. Each participant completed two procedures: phrase-to-force translation and force-to-phrase translation. Both procedures used a consistent reference frame as in (Section III-A).
@@ V-C1 Phrase-to-Force
Participants were sequentially presented with 42 distinct phrases (Table  II illustrates an example sequence of prompts.). Each phrase was generated by randomly combining a direction word (e.g., “left”, “forward-down”) with an optional modifier word (e.g., “quickly”, “smoothly”) from the vocabulary described in Section  III-C. The first six phrases were always basic directions, while the remaining 36 incorporated modifiers and compound directions. For each phrase, participants gripped the robot’s end-effector and demonstrated what they felt was an appropriate motion matching the description. The system recorded force measurements for each demonstration, typically lasting 2-4 seconds. A user trial demonstrating this is shown in Figure  7(a).
@@ V-C2 Force-to-Phrase
In this procedure, participants gripped the robot’s end-effector while the manipulator executed randomly generated force trajectories. Each motion was created by combining three randomized parameters: a primary direction vector, a force magnitude between 0.5 and 15 Newtons, and a duration ranging from 1 to 4 seconds. After experiencing each motion, participants constructed descriptive phrases using the same vocabulary (Section  III-C) from the phrase-to-force procedure - selecting direction and modifier words they felt best characterized the force profile they had just experienced. A user participant trial demonstrating this is shown in Fig.  7(b). Complete details on the dataset with all 10 participants are in appendix  [^4] Online Appendix showcasing the data collected from 10 user participants for phrase-to-force and force-to-phrase https://shared-language-force-embedding.github.io/data-collection/.
@@ VI Evaluation
    Impact of Phrase Representation. How does the choice of phrase representation affect model performance? Specifically, do richer semantic embeddings (via pretrained GloVe embeddings) enhance the model’s ability to associate forces with language compared to binary phrase vectors from the Minimal Viable Vocabulary?
To investigate these questions, we evaluate two variants of our dual autoencoder (DAE) framework. The first variant, denoted $DAE_{B}$, utilizes the binary phrase vector representation of phrases (described in Section  III-C2). The second variant, $DAE_{G}$, employs the GloVe embedding representation (described in Section  III-C2. By comparing results across these two models, we assess the influence of language embedding choices on overall performance and generalization.
@@ VI-A1 $SVM\_KNN$
As a baseline, we use Support Vector Machines (SVM) to map force signals to phrases and K-Nearest Neighbors (KNN) to map phrases back to force profiles. This approach is limited to our Minimal Viable Vocabulary since SVMs cannot generate continuous GloVe embeddings.
@@ VI-A2 $DMLP_{B}$
VI-A3 $DMLP_{G}$
$DMLP_{G}$ is analogous to $DMLP_{B}$ except it uses the GloVe embedding representation of phrases (Section  III-C2). Instead of two binary phrase vector outputs, the first MLP maps the force profile to a 150-dimensional concatenation of GloVe word embeddings, and the second MLP performs the inverse mapping. Each MLP is again optimized with a reconstruction-style loss suited to its respective output space. This direct mapping baseline allows us to isolate the effect of introducing richer semantic information via GloVe embeddings, independent of a shared latent representation.
@@ VI-C Experiments
(a) Force profile predictions for “quickly up” representing an in-distribution phrase
(b) Force profile predictions for “greatly forward” testing generalization for the out-of-distribution phrase.
(c) Force profile predictions for the out-of-distribution direction “backward up”.
Figure 8: Visualization of predicted force profiles in X, Y, and Z axes (rows) from our model variants ($DAE_{G}$ and $DAE_{B}$) compared to ground truth models ($SVM\_KNN$, $MLP_{B}$ and $MLP_{G}$) for experiments in Section VI-C. Results for all the possible combination of phrases on each baseline model can be found in our online appendix  [^6] Online Appendix showcasing the results for all the combinations of phrases for each baseline models https://shared-language-force-embedding.github.io/results/
Model Modifier Direction
Reference quickly up
SVM_KNN - up
$MLP_{B}$ quickly up
$MLP_{G}$ meantime up
$DAE_{B}$ sharply up
$DAE_{G}$ quickly up
(a) In-distribution phrase predictions.
Model Modifier Direction
Reference greatly forward
SVM_KNN harshly forward
$MLP_{B}$ forward
$MLP_{G}$ likewise forward
$DAE_{B}$ significantly forward
$DAE_{G}$ significantly forward
(b) Out-of-distribution modifier predictions.
Model Direction 1 Direction 2
Reference backward up
SVM_KNN backward -
$MLP_{B}$ backward -
$MLP_{G}$ backward -
$DAE_{B}$ backward -
$DAE_{G}$ backward -
(c) Out-of-distribution direction predictions.
Figure 9: Example phrase predictions from each experiment. Phrases are predicted from the corresponding ground truth force profiles shown in Fig. 6
@@ VI-D Results and Analysis
We present our experimental results addressing three research questions (see VI for complete descriptions):
@@
    What is the impact of the phrase representation (binary vs. GloVe) on performance?
  1. 1.
    In-Distribution Results When trained and tested on the same distribution of force–phrase pairs, all four neural approaches ($\text{DMLP}_{B}$, $\text{DMLP}_{G}$, $\text{DAE}_{B}$, $\text{DAE}_{G}$) performed substantially better than SVM_KNN in terms of reproducing the force profile (FPAcc) and capturing the aggregate direction (FDAcc).
@@
    Additionally, $\text{DAE}_{B}$ and $\text{DMLP}_{B}$ both outperformed $\text{DAE}_{G}$ and $\text{DMLP}_{G}$ on in-distribution data. This suggests that the simpler binary phrase vectors may be sufficiently discriminative—and perhaps easier to learn—when the training distribution covers the same words. This observation partially answers our third question by illustrating that the binary representation can be advantageous for in-distribution translation.
  2. 2.
    Out-of-Distribution Modifiers For unseen modifiers, $\text{DAE}_{G}$ achieves the best force profile reconstruction (FPAcc: 5.41) while $\text{DAE}_{B}$ maintains the highest phrase similarity scores (PhraseSim: 0.68). This indicates that the DAE models retain semantic representation of adverbs even if they never explicitly see them during training. We observe that the DAE models generalize more effectively to unseen modifiers than MLPs. GloVe embeddings helps in reconstructing the force trajectory (FPAcc), but the binary phrase vector representation still achieves strong direction and phrase-level fidelity. This addresses our second question: the dual autoencoders can indeed handle certain unseen modifiers better than simpler baselines, although each representation has trade-offs.
  3. 3.
    Out-of-Distribution Directions Both $\text{DAE}_{G}$ and $\text{DMLP}_{G}$ show superior force profile reconstruction (FPAcc) and directional accuracy (FDAcc) compared to binary-based models for unseen directions, suggesting that GloVe’s continuous semantics facilitate better extrapolation of physical orientation and overall impulse direction for unseen directional words.
    Conversely, the binary-vector based models achieved higher scores in ModSim, DirSim and PhraseSim, indicating that although $\text{DAE}_{G}$ and $\text{DMLP}_{G}$ produce more faithful physical force reconstructions, their textual outputs are less aligned with the precise wording of the missing direction. This addresses our third question: the GloVe representation demonstrates stronger zero-shot direction encoding on the force side, but experiences a penalty in reproducing the textual labels for directions.
Key Findings These results demonstrate that:
  • The dual autoencoder architecture effectively enables bidirectional translation, with DAE variants showing 20-30% improvement over baselines;
  • Models can generalize to unseen inputs, with GloVe variants showing particular strength ( 25% improvement) in force reconstruction;
  • Representation choice presents a trade-off between force reconstruction accuracy (favored by GloVe) and linguistic precision (favored by binary encoding). The choice between binary phrase vectors and GloVe embeddings representations should be guided by application requirements: GloVe for better force reconstruction and generalization, or binary phrase vector representation for precise linguistic mapping and in-distribution performance.

### v2

# Cross-modality Force and Language Embeddings for Natural Human-Robot Communication

@@ I Introduction
As stipulated in Asimov’s laws, robots must perform tasks based on human instruction. A central challenge in robotics has been developing effective ways for humans to communicate and give instructions to robots.
Figure 1: Physical therapist from Spaulding Rehabilitation Hospital is seen demonstrating ’hamstring curl’ therapy on the patient with neurological injuries  [^2] Detailed therapy sessions can be seen in our online appendix: https://shared-language-force-embedding.github.io/therapy-sessions/. She is instructing a patient to move ’gently forward’ while providing an assistive force.
Early roboticists attempted to develop robot languages, e.g. VAL [19], to describe desired robot actions and tasks. However, despite some success, it became apparent that certain motions and behaviors are difficult, inefficient, or impossible to describe through language alone. This limitation is particularly evident in tasks requiring environmental contact and force/compliance control, where forces and moments are not directly visible. Such manipulative tasks, now termed contact-rich manipulation [20, 7], often involve subconscious knowledge that humans find difficult to articulate through any form of language. To address this challenge, roboticists developed alternative approaches that bypass the need to translate subconscious skills into language. These approaches include teaching by demonstration [3], programming by demonstration [6], skill acquisition [17] and, more recently, imitation learning [25].
@@
    A framework for cross-modality embedding of force profiles and words  [^3] Demonstration: https://shared-language-force-embedding.github.io/demo/, enabling translation between physical force curves and natural language descriptions;
@@ III-C1 Minimal Viable Vocabulary
III-C2 Extended SBERT Vocabulary
We leverage SBERT (Sentence-BERT) embeddings [23], a contextual language model that produces semantically meaningful representations. This representation allows for encoding beyond our minimal viable vocabulary, as the SBERT model has been pretrained on diverse textual data. When processing a phrase, we encode the complete text rather than individual words, resulting in a 768-dimensional vector that preserves the semantic meaning of the entire expression (Fig. 4). This approach enables our system to handle open-vocabulary inputs while maintaining the shared structure of our force-language embedding space.
Figure 3: Conceptual illustration of the desired properties of the cross-modality latent space. A pair of corresponding force profile and phrase should be located near each other, measured by a distance metric such as cosine similarity. However, force profiles and phrases that do not correspond should be positioned far away. This demonstrates that similar inputs would be close together and dissimilar inputs would be far apart in the latent space.
Figure 4: Conceptual illustration of the process of matching an arbitrary text input with the most semantically similar phrase using SBERT embeddings.
@@ III-D Cross-Modality Embedding (Shared Latent Space)
Figure 6: Architecture diagram of the cross-modality dual autoencoder that represents phrases as a $768\times 1$ embedding generated by S-BERT ([23]). The phrase embedding input is then passed into the phrase autoencoder, which encodes it into the shared latent space. There the $16\times 1$ embedding can take 2 paths: either be decoded back into a phrase or be translated into a force profile by using the force profile decoder. Force profile inputs are first preprocessed to extract meaningful features for the force profile autoencoder to digest. They are then encoded into the shared latent space. Like phrase inputs, they can either be decoded back into a force profile or be translated into a phrase by using the phrase decoder.
@@ V-A2 Phrase Input
Since neural networks cannot directly process raw text, each phrase must be converted into a fixed-size vector representation that preserves its semantics. We consider two embedding approaches: Binary Phrase Vectors from the Minimal Viable Vocabulary (Section  III-C1), and contextual embeddings from SBERT (Section  III-C2).
Binary Phrase Vectors
In our minimal viable vocabulary, each phrase is composed of exactly two words: a modifier and a direction. Each word is represented as a one-hot encoded vector, with the modifier requiring 31 dimensions and the direction requiring 31 dimensions (including an ”empty” position for each). These binary vectors are then concatenated to yield a 62-dimensional phrase representation. This discrete encoding provides a direct mapping between the linguistic components and physical force characteristics.
SBERT Embeddings
To handle arbitrary verbal instructions beyond our fixed vocabulary, we implement a mapping mechanism using SBERT that identifies the most semantically similar phrase in our minimal viable vocabulary. For any input text $x$, we find the matching MVV phrase $p^{*}$ using:
$p^{*}=\begin{cases}\arg\max_{p\in P}s(E(p),E(x))\;&\text{if }s(E(p),E(x))>\sigma\\ \text{``"}\;&\text{otherwise}\end{cases}$ (5)
where $P$ is the set of MVV phrases, $E$ is the SBERT encoder, $s$ is cosine similarity, and $\sigma$ is a threshold parameter. This approach enables our system to accept open-vocabulary input while leveraging our trained force-language mappings.
@@ V-A3 Dual Autoencoders
Our framework employs two autoencoders [15] —one for force profiles and one for phrases—to map inputs from distinct modalities into a shared latent space. For the phrase autoencoder, we utilize SBERT’s 768-dimensional embeddings or the Binary Phrase Vector (62-dimensional embedding) as input, which are then processed through several layers before reaching the 16-dimensional bottleneck layer. This architecture enables the model to compress the semantic information into a compact representation that aligns with force profiles in latent space.
@@ V-B Multitask Learning
$\mathcal{L}=k_{r}\mathcal{L}_{r}+k_{z}\mathcal{L}_{c}+k_{t}\mathcal{L}_{t},$ (6)
@@ V-B1 Reconstruction Loss ($\mathcal{L}_{r}$)
It measures how accurately each autoencoder (force and phrase) reproduces its own input from the latent vector. For force profiles, we use mean squared error; for phrases, the reconstruction metric depends on the chosen representation, with cross-entropy for binary phrase vectors and mean squared error for S-BERT embeddings.
For phrase inputs $x_{p}$ with encoder $E_{p}$ and decoder $D_{p}$, the reconstruction loss varies:
$\mathcal{L}_{r,p}(x_{p})=\begin{cases}H(x_{p},\text{softmax}(D_{p}(E_{p}(x_{p}))))&\text{binary}\\ \|x_{p}-D_{p}(E_{p}(x_{p}))\|^{2}&\text{SBERT}\end{cases}$ (7)
where $H$ is cross-entropy loss.
@@ V-B2 Contrastive Loss ($\mathcal{L}_{c}$)
The contrastive loss for a batch of $n$ force-phrase pairs $(z_{f}^{i},z_{p}^{i})$ in the shared latent space is:
$\mathcal{L}_{c}=\sum_{i=1}^{n}\|z_{f}^{i}-z_{p}^{i}\|^{2}-\lambda\sum_{i=1}^{n}\sum_{j\neq i}^{n}\max(0,m-\|z_{f}^{i}-z_{p}^{j}\|^{2})$ (8)
where $\lambda$ controls the negative pair weighting and $m$ is the margin parameter.
@@ V-B3 Translation Loss ($\mathcal{L}_{t}$)
The translation loss for corresponding force profile and phrase inputs $x_{f}$ and $x_{p}$ is:
$\mathcal{L}_{t}(x_{f},x_{p})=\text{err}(x_{f},D_{f}(E_{p}(x_{p})))+\text{err}(x_{p},D_{p}(E_{f}(x_{f})))$ (9)
where $\text{err}(\cdot,\cdot)$ is the appropriate reconstruction error metric for each modality.
@@ V-C Data Collection
To train and evaluate the cross force-language embedding, we collected data from 10 participants (6 males, 4 females) aged 21-32 years (mean=26.4) who interacted with a UR10 manipulator arm. To address individual variations in force profile application, we employed min-max normalization and relative pattern analysis across participants. This normalization procedure helps account for differences in strength and movement patterns between individuals, allowing the model to focus on the semantic relationship between language and force trajectories rather than absolute force magnitudes. Each participant completed two procedures: phrase-to-force translation and force-to-phrase translation. Both procedures used a consistent reference frame as in (Section III-A). The study was approved by the Institutional Review Board (IRB Protocol #2212000845R001) to ensure ethical human subject research guidelines were followed.
@@ V-C1 Phrase-to-Force
Participants were sequentially presented with 42 distinct phrases (Table  II illustrates an example sequence of prompts.). Each phrase was generated by randomly combining a direction word (e.g., “left”, “forward-down”) with an optional modifier word (e.g., “quickly”, “smoothly”) from the vocabulary described in Section  III-C. The first six phrases were always basic directions, while the remaining 36 incorporated modifiers and compound directions. For each phrase, participants gripped the robot’s end-effector and demonstrated what they felt was an appropriate motion matching the description. The system recorded force measurements for each demonstration, typically lasting 2-4 seconds. A user trial demonstrating this is shown in Figure  8(a).
@@ V-C2 Force-to-Phrase
In this procedure, participants gripped the robot’s end-effector while the manipulator executed randomly generated force trajectories. Each motion was created by combining three randomized parameters: a primary direction vector, a force magnitude between 0.5 and 15 Newtons, and a duration ranging from 1 to 4 seconds. After experiencing each motion, participants constructed descriptive phrases using the same vocabulary (Section  III-C) from the phrase-to-force procedure - selecting direction and modifier words they felt best characterized the force profile they had just experienced. A user participant trial demonstrating this is shown in Fig.  8(b). Complete details on the dataset with all 10 participants are in appendix  [^5] Online Appendix showcasing the data collected from 10 user participants for phrase-to-force and force-to-phrase https://shared-language-force-embedding.github.io/data-collection/.
@@ VI Evaluation
    Impact of Phrase Representation. How does the choice of phrase representation affect model performance? Specifically, do richer semantic embeddings (via pretrained S-BERT embeddings) enhance the model’s ability to associate forces with language compared to binary phrase vectors from the Minimal Viable Vocabulary?
To investigate these questions, we evaluate two variants of our dual autoencoder (DAE) framework. The first variant, denoted $DAE_{B}$, utilizes the binary phrase vector representation of phrases (described in Section  III-C2). The second variant, $DAE_{S}$, employs the SBERT embedding representation (described in Section  III-C2. By comparing results across these two models, we assess the influence of language embedding choices on overall performance and generalization.
@@ VI-A1 $SVM\_KNN$
As a baseline, we use Support Vector Machines (SVM) to map force signals to phrases and K-Nearest Neighbors (KNN) to map phrases back to force profiles. This approach is limited to our Minimal Viable Vocabulary since SVMs cannot generate continuous S-BERT embeddings.
@@ VI-A2 $DMLP_{B}$
VI-A3 $DMLP_{S}$
$DMLP_{S}$ is analogous to $DMLP_{B}$ except it uses the S-BERT embedding representation of phrases. Instead of two binary phrase vector outputs, the first MLP maps the force profile to a 150-dimensional concatenation of S-BERT word embeddings, and the second MLP performs the inverse mapping. Each MLP is again optimized with a reconstruction-style loss suited to its respective output space. This direct mapping baseline allows us to isolate the effect of introducing richer semantic information via S-BERT embeddings, independent of a shared latent representation.
@@ VI-C Experiments
Figure 9: Comparative force profiles for the open vocabulary phrase “Ease back and up a bit more” across different models. Solid lines represent ground truth reference forces, dashed lines show predictions from our baseline DMLPS model, and dotted lines show predictions from our proposed DAES model. The negative $Y$ component (green) correctly captures the “back” direction with both models showing similar trajectory shapes but different peak magnitudes. The positive $Z$ component (blue) represents the “up” direction, with DAES more closely matching the reference profile shape though with a smoother decay. The $X$ component (red) shows minimal activation as expected since the phrase doesn’t specify rightward movement. This visualization demonstrates how our models interpret complex verbal instructions containing both direction and intensity modifiers. For complete set of force curve examples and phrases, see our online appendix. [^5] Online Appendix showcasing the results for all the combinations of phrases for each baseline models https://shared-language-force-embedding.github.io/results/
@@ VI-D Results and Analysis
We present the evaluation outcomes for the three experimental settings: in-distribution, out-of-distribution (OOD) modifiers, and out-of-distribution directions. Recall our original research questions:
@@
    What is the impact of the phrase representation (binary vs. S-BERT) on performance?
Our analysis of the experimental results addresses each of these questions:
In-Distribution Results
When trained and tested on the same distribution of force–phrase pairs, all four neural approaches ($\text{DMLP}_{B}$, $\text{DMLP}_{S}$, $\text{DAE}_{B}$, $\text{DAE}_{S}$) performed substantially better than SVM_KNN in terms of reproducing the force profile (FPAcc) and capturing the aggregate direction (FDAcc).
@@
Figure 11: Visualization of predicted force profiles using our SBERT-based similarity mapping for open vocabulary inputs. (a) For the directional command “Let’s go upward and to the right,” the system generates appropriate force profiles with dominant components in the z-axis (up) and x-axis (right). (b) The more abstract instruction “I want you to ease ahead over time” is mapped to a gradually increasing y-axis (forward) force with appropriate duration. (c) For the non-force related phrase “I like apples,” the system produces minimal force output, demonstrating the ability to distinguish between force commands and unrelated statements. Force components are shown in Newtons (N) over a 4-second duration.
@@
Additionally, $\text{DAE}_{B}$ and $\text{DMLP}_{B}$ both outperformed $\text{DAE}_{S}$ and $\text{DMLP}_{S}$ on in-distribution data. This suggests that the simpler binary phrase vectors may be sufficiently discriminative—and perhaps easier to learn—when the training distribution covers the same words. This observation partially answers our third question by illustrating that the binary representation can be advantageous for in-distribution translation.
Out-of-Distribution Modifiers
For unseen modifiers, $\text{DAE}_{S}$ achieves the best force profile reconstruction (FPAcc: 5.41) while $\text{DAE}_{B}$ maintains the highest phrase similarity scores (PhraseSim: 0.68). This indicates that the DAE models retain semantic representation of adverbs even if they never explicitly see them during training. We observe that the DAE models generalize more effectively to unseen modifiers than MLPs. S-BERT embeddings helps in reconstructing the force trajectory (FPAcc), but the binary phrase vector representation still achieves strong direction and phrase-level fidelity. This addresses our second question: the dual autoencoders can indeed handle certain unseen modifiers better than simpler baselines, although each representation has trade-offs.
Out-of-Distribution Directions
Both $\text{DAE}_{S}$ and $\text{DMLP}_{S}$ show superior force profile reconstruction (FPAcc) and directional accuracy (FDAcc) compared to binary-based models for unseen directions, suggesting that S-BERT’s continuous semantics facilitate better extrapolation of physical orientation and overall impulse direction for unseen directional words. Conversely, the binary-vector based models achieved higher scores in ModSim, DirSim and PhraseSim, indicating that although $\text{DAE}_{S}$ and $\text{DMLP}_{S}$ produce more faithful physical force reconstructions, their textual outputs are less aligned with the precise wording of the missing direction. Hence, the S-BERT representation demonstrates stronger zero-shot direction encoding on the force side, but experiences a penalty in accurately reproducing the textual labels for directions.
Summary Fig 10 shows the comprehensive performance comparison of our method across in-distribution testing and generalization scenarios. These results demonstrate that: (1) Dual autoencoder architecture effectively enables bidirectional translation, with DAE variants showing 20-30% improvement over baselines; (2) Models can generalize to unseen inputs, with S-BERT variants showing particular strength ( 25% improvement) in force reconstruction; (3) Representation choice presents a clear trade-off between force reconstruction accuracy (favored by S-BERT) and linguistic precision (favored by binary encoding). The choice between binary phrase vectors and S-BERT embeddings representations should be guided by application requirements: S-BERT for better force reconstruction and generalization, or binary phrase vector representation for precise linguistic mapping and in-distribution performance.